OpenStudy (anonymous):

which ordered pair is a solution of the equation y=1/2 +3

OpenStudy (mertsj):

Did you mean y = 1/2x +3? And did you mean to list some ordered pairs?

OpenStudy (anonymous):

If so, just sub them in, if LS = RS, it's the right pair.

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

possible answers 5,-4 4,-1 4,1 8,1

OpenStudy (mertsj):

Let's try them one at a time until we find one that makes a true statement.

OpenStudy (mertsj):

\[-4=\frac{1}{2}(5)+3=\frac{5}{2}+3\] That one is not true. Let's try the next one.

OpenStudy (mertsj):

\[-1=\frac{1}{2}(4)+3=2+3=5\] Since -1 is not equal to 5, that doesn't work either. Let's try the next one.

OpenStudy (mertsj):

\[1=\frac{1}{2}(4)+3=2+3=5\] That's not true either. Let's try the next one.

OpenStudy (mertsj):

\[1=\frac{1}{2}(8)+3=7\] That's not true either. None of the ordered pairs you posted make the sentence true so apparently you have posted something incorrectly.